OpenStudy (anonymous):

I fan and give medals. :) A prism has a volume of 16 cm3. What is the volume of the prism in cubic millimeters? 160,000 mm3 16,000 mm3 160 mm3 1,600 mm3

OpenStudy (anonymous):

Is it a rectangular prism?

OpenStudy (anonymous):

@Isaiah.Feynman could you help?

OpenStudy (anonymous):

IDK it just says a prism

OpenStudy (anonymous):

A centimeter has 100 meters A millimeter has 1000 meters This means that 16 cm3 is equal to 160 millimeters.

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

Yes it is XD I just confirmed it. Does that make sense to you Beth?

OpenStudy (anonymous):

No not really....

OpenStudy (anonymous):

16cm3 = 160 mm3

OpenStudy (anonymous):

\[16cm^3=16cm^3 * (\frac{10mm}{1cm})^3=16cm^3*\frac{1000mm^3}{1cm^3}=16000mm^3\]

OpenStudy (anonymous):

That was wrong

OpenStudy (anonymous):

16cm3 is 16 x 100 = 1600 meters3. Do i have you so far?

OpenStudy (anonymous):

1 cm3 = (1 cm) × (1 cm) × (1 cm) = 10 mm × 10 mm × 10 mm = 1,000 mm3 So, 16 cm3 = 16 cm3 × 1,000 mm3/cm3 = 16,000 mm3

OpenStudy (anonymous):

That's what I put

OpenStudy (anonymous):

To do this easier, just search google "cm to mm" and a calculator comes and and you enter the cm/ or/ mm. It depends on if you want to find out cm or mm. If you have to do the hard way here is the equation: \[mm = \frac{ cm }{ 0.10000 }\]

OpenStudy (anonymous):

Ah because this is volume. If you are looking for volume then you need to move more exponents.

OpenStudy (anonymous):

Thanks for the help though!:)